OpenStudy (anonymous):

how to find the limit of the f(x) sin(2x)/4x as it approaches 0 without a calculator

OpenStudy (anonymous):

ok well, do you know the limit of sinx/x as it approaches 0?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

you will need to use a trig identity

OpenStudy (anonymous):

oh nevermind no you won't. you just have to plug in 0

OpenStudy (anonymous):

right, so you have 2x instead of x, you want to get a 2x on the bottom, how would you do that with 4x?

OpenStudy (anonymous):

plugging in 0 would get you 0/0 which is no bueno

OpenStudy (anonymous):

ok so then yes you will need an identity

OpenStudy (anonymous):

you want to get 2x on the bottom, so you would have to divide by 2

OpenStudy (anonymous):

and you can take \[\frac{ 1 }{ 4 }\] as a constant

OpenStudy (anonymous):

right, you would factor out a 2

OpenStudy (anonymous):

you could take 1/4 out as a constant, but then you'd still have sin(2x)/x in your limit. Then you would need to multiple in 2/2 to get 2sin(2x)/2x

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\frac{ \sin(2x) }{ 2x} * \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

then you can use your limit identity, but you'd get the same answer.

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

now, what's the \[\lim_{x \rightarrow 0}\frac{ \sin(2x) }{ 2x }\]

OpenStudy (anonymous):

1 still, right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

all you're left with is 1/2 and that's your answer. :)

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

No worries. :)